OpenStudy (anonymous):

A city’s current population is 1,000,000 people. It is growing at a rate of 3.5% per year. The equation P = 1,000,000(1.035)x models the city’s population growth where x is the number of years from the current year. In approximately how many years will the population be 1,200,000? Round to the nearest tenth.

OpenStudy (aaronq):

You use the equation given. P = 1,000,000(1.035)x P=is the population after x years so, P=1,200,000=1,000,000(1.035)x solve for x

OpenStudy (anonymous):

I got 5.3

OpenStudy (aaronq):

i got 1.2

OpenStudy (anonymous):

Its not 1 I thought it was too but its wronggg

OpenStudy (aaronq):

1,200,000/1,000,000=1.035x 1.2=1.035x 1.2/1.035=x x=1.1594202898550725 they wanted rounded to the nearest tenth x=1.2

OpenStudy (anonymous):

that's rounded to 1 ? I tried 1 it was wrong

OpenStudy (aaronq):

round to 1.2, tenth decimal place

OpenStudy (anonymous):

its 0

OpenStudy (anonymous):

that can't be right

OpenStudy (aaronq):

zero? lol it can't be zero, thats nonsense.

OpenStudy (anonymous):

zero!!

OpenStudy (anonymous):

That's what that said

OpenStudy (anonymous):

\[1.2=(1.035)^x\iff x=\frac{\ln(1.2)}{\ln(1.035)}\]

OpenStudy (anonymous):

1.1594

OpenStudy (anonymous):

and a calculator the confusion was, the \(x\) is in the exponent

OpenStudy (anonymous):

that's what I got ?

OpenStudy (anonymous):

your calculator is not working right http://www.wolframalpha.com/input/?i=ln%281.2%29%2Fln%281.035%29

OpenStudy (anonymous):

soo its 10 ?

OpenStudy (anonymous):

@satellite73